# JailbreakLens: Interpreting Jailbreak Mechanism in the Lens of Representation and Circuit

### v1

@@ Abstract
Despite the outstanding performance of Large language models (LLMs) in diverse tasks, they are vulnerable to jailbreak attacks, wherein adversarial prompts are crafted to bypass their security mechanisms and elicit unexpected responses. Although jailbreak attacks are prevalent, the understanding of their underlying mechanisms remains limited. Recent studies have explain typical jailbreaking behavior (e.g., the degree to which the model refuses to respond) of LLMs by analyzing the representation shifts in their latent space caused by jailbreak prompts or identifying key neurons that contribute to the success of these attacks. However, these studies neither explore diverse jailbreak patterns nor provide a fine-grained explanation from the failure of circuit to the changes of representational, leaving significant gaps in uncovering the jailbreak mechanism. In this paper, we propose JailbreakLens, an interpretation framework that analyzes jailbreak mechanisms from both representation (which reveals how jailbreaks alter the model’s harmfulness perception) and circuit perspectives (which uncovers the causes of these deceptions by identifying key circuits contributing to the vulnerability), tracking their evolution throughout the entire response generation process. We then conduct an in-depth evaluation of jailbreak behavior on four mainstream LLMs under seven jailbreak strategies. Our evaluation finds that jailbreak prompts amplify components that reinforce affirmative responses while suppressing those that produce refusal. Although this manipulation shifts model representations toward safe clusters to deceive the LLM, leading it to provide detailed responses instead of refusals, it still produce abnormal activation which can be caught in the circuit analysis. Additionally, we uncover unique distinctive characteristics of each jailbreak method and examine how model scale and fine-tuning influence vulnerability to these attacks.
1 Introduction
@@
Large language models (LLMs) [1, 2] have revolutionized various fields with powerful capabilities. While LLMs acquire extensive knowledge from massive corpora during pre-training, they also inevitably embed harmful knowledge, which poisons LLMs to generate harmful responses that may violate ethics. To mitigate such harmful outputs, LLMs are further aligned via reinforcement learning to develop security mechanisms[3, 4, 5] into them.
@@
However, flaws in the alignment process persist [6, 7, 8, 9, 10], as evidenced by thousands of jailbreaks, that attackers bypass the safety mechanisms by using carefully crafted prompts. Prevalent jailbreak methods can be generally categorized into five types [11], i.e., gradient-based attacks [6], evolutionary-based attacks [7, 12], demonstration-based attacks [8], rule-based attacks [9, 13], and multi-agent-based attacks [10]. These attacks have led to increasing security incidents, raising serious concerns about the safety of LLM responses. A tragic example [14] is the suicide of a 14-year-old boy suffering from depression who had become heavily dependent on his AI companion.
@@
To understand the failure modes of safety alignment in real-world scenarios, interpreting the internal states of LLMs under jailbreak threats has become crucial for developing trustworthy super intelligence. Existing research [15, 16, 17, 18] has explored how LLMs process and represent harmful prompts, showing that jailbreak prompts can be disguised within benign clusters, making it hard for well-aligned models to differentiate between safe and harmful prompts. While representation analysis reveals the shifts in latent space, it cannot trace the specific internal component to uncover the causes of these changes. To address this, some studies [19, 20] focus on circuit analysis, particularly examining the MLP layers’ neurons responsible for safety behaviors. Circuit analysis, on the other hand, usually focuses on a limited set of modules, such as certain neurons. Since jailbreak often involves interactions across multiple layers, the effects of which are amplified deeper into the model, it is difficult to capture the impact on individual layers and overall behavior by analyzing only a limited set of circuits.
While existing works have taken pioneering steps in exploring jailbreak mechanisms, we identify that there are still inherently several gaps. (1). Limitations in Jailbreak Types. Existing studies mainly focus on limited types of classic jailbreak methods, such as GCG and PAIR, etc., and thus fail to cover diverse state-of-the-art jailbreak strategies that exhibit novel attack features. (2). Insufficient Integration Between Representation and Circuits Analysis. Representation analysis reveals how jailbreaks deceive the model’s harmfulness perception, preventing it from recognizing malicious intent within the prompts, while circuit analysis uncovers the cause of this deception by identifying key circuits contributing to the vulnerability. Combining both can better reveal how jailbreak attacks penetrate the model’s behaviors layer by layer and identify the underlying circuits involved. (3). Lack of Dynamic Analysis. Most studies focus on static analysis to examine the first token generation of jailbreak responses, failing to capture the fine-grained jailbreak behaviors throughout the sentence-level generation. Such lack leads to inaccurate interpretations. For example, a model may initially respond safely, but gradually become affected by a jailbreak prompt and generate harmful content.
@@
Figure 1: An overview of JailbreakLens, a framework that interprets LLM jailbreak behavior via capturing representations and circuits. Given an LLM, JailbreakLens first processes a jailbreak prompt to observe intermediate representation and model activations. In Representation Probing Stage, we predict the safety score of representation with a pre-trained safety probing classifier to detect whether the model identify the harmfulness within the jailbreak prompts. In Model Behavior Circuit stage, we map the safety direction vector identified by the probing classifier into token space to obtain two token sets (i.e., affirmation tokens and refusal tokens). By attributing activation differences between the two sets, we identify key circuits where activation changes significantly and investigate how their behaviors are impacted by the jailbreak prompt. Finally, we integrate sentence-level safety scores with key circuit activations to derive the final interpretation results.
To address these gaps, we propose JailbreakLens, a dual-perspective interpretation framework that can provide a fine-grained accurate interpretation of LLM jailbreaks from both representation level and circuit perspectives. We first capture the changes caused by the jailbreak prompt at the representation level, then dig deeper into the functional reasons at circuit level. Finally we combine the two to provide a comprehensive explanation linking representation shifts to the circuit behaviors. To examine how jailbreak prompt obscure harmfulness perception within the representation space, we predict the safety score of representation with a pre-trained probing classifier to detect whether the model identifies the harmfulness within the jailbreak prompt. At the circuit level, we aim to locate key components contributing to safety and investigate how jailbreak prompt disrupt their behaviors. Specifically, we map the safety direction vector identified by the probing classifier into token space to obtain two token sets (i.e., affirmation tokens and refusal tokens), since the direction vector provides a linear representation of how safety-related features are encoded, allowing us to decode the safety perception into token-level semantics. By attributing the model to the differences between the two sets, we identify key circuits where activation changes significantly. Finally, we integrate the sentence-level safety scores and the key circuit activations to derive the final interpretation of the jailbreak prompt. Our evaluation finds that jailbreak prompts amplify components reinforcing affirmative responses while suppressing those that produce refusal responses, leading the model representations shift toward safe clusters to deceive LLM, however, it still produce abnormal activation which can be caught in circuit analysis.
@@
  • In-depth Jailbreak Mechanism Analysis. We propose JailbreakLens, a interpretation framework that examines how jailbreak attacks break LLMs’ safeguard from the perspectives of internal representations and circuits, addressing the main gaps of jailbreak mechanism interpretation in current research.
  • Comprehensive Study of Jailbreak Strategies. We investigate seven jailbreak methods across five major categories, on four mainstream LLMs. Extensive experiments provide insights into how various jailbreak types exploit model vulnerabilities across different architectures and scales.
  • Extensive Findings on Jailbreak Mechanism. We identify many findings with our proposed interpretation framework. For instance, jailbreak prompts amplify components reinforcing affirmative responses by 200 times more than harmful prompts, while suppressing those that produce refusal responses by 45% on average. Although, this manipulation shifts model representations toward safe clusters to deceive LLM, leading it to provide detailed responses instead of refusals, it still produce abnormal activation which can be caught in circuit analysis. Moreover, we identify that increasing parameter scale does not significantly improve the model’s alignment generalization capacity, and instruction tuning has minimal impact on the key circuit location of safeguard.
2 Related Works
@@ 2.1 Jailbreak Attacks on LLM
LLM jailbreak denotes that attackers strategically manipulate the input prompts with the intent to bypass the LLM’s safeguards, exploiting vulnerabilities across dimensions [21] like hallucinations, violence, bias, and mental health risks. Following JailbreakZoo [11], we broadly categorize the jailbreak methods into five main types based on the prompt generation strategy, i.e., gradient-based [6], evolutionary-based [7, 12], demonstration-based [8], rule-based [13, 9], and multi-agent-based [10] jailbreaks.
@@
Gradient-based jailbreaks exploit the gradients of the model to optimize the inputs. For example, Greedy Coordinate Gradient (GCG) [6] generates a suffix that can elicit harmful responses when attached to a broad spectrum of queries directed at a targeted LLM. The suffix is calculated by greedy search from random initialization to maximize the likelihood that the model produces an affirmative response.
@@
Evolutionary-based jailbreaks generate adversarial prompts via genetic algorithms and evolutionary strategies. For example, Yu et al. [7] propose GPTFuzzer that incorporates the concept of mutation, initiating with human-crafted templates as the foundational seeds, and subsequently mutating these seeds to generate novel templates. Moreover, Liu et al. [12] introduce Autodan which can automatically generate stealthy jailbreak prompts by the carefully designed hierarchical genetic algorithm.
@@
Demonstration-based jailbreaks use fixed, carefully crafted prompts to elicit specific responses from a language model, without adapting the prompt to different queries. For example, DeepInception [8] leverages the personification capabilities of LLMs by creating nested scene prompts that engage the model in complex, multi-layered contexts, subtly bypassing the safety guardrails.
@@
Rule-based jailbreaks decompose and redirect malicious prompts through predefined rules to evade detection. ReNeLLM [13] firstly rewrites the initial prompts with a series of rewriting operations without altering its core semantics and then nest the rewritten prompt within universal task scenarios such as table filling. CodeChameleon [9] elude the intent safety guardrail by reformulating tasks into a code completion format, enabling users to encrypt queries using personalized encryption functions.
@@
Multi-agent-based jailbreaks depend on the cooperation of multiple LLMs to iteratively refine and enhance jailbreak prompts. Methods such as PAIR [10] use the collaboration of a red team LLM to optimize and improve the effectiveness of jailbreak prompts.
@@
To investigate diverse jailbreak strategies, we select 1-2 specific methods from each category, addressing the limited exploration of jailbreak categories in previous work on jailbreak mechanism explanations.
2.2 LLM Mechanism Interpretability
Research on LLM mechanism interpretability aims to provide transparency and interpretability on LLM’s behaviors. Attribution methods include behavior localization through input attribution [22, 23], which localizes behaviors by tracing model responses back to specific inputs, and model component attribution [24, 25], which identifies influential model components responsible for specific outputs. Causal interventions, like activation patching[26, 27], reveal critical components by altering internal structures and observing the effects on predictions. The linear representation hypothesis[28] posits that features are encoded in subspaces, with techniques like probing[29, 30] and sparse auto-encoders[31, 32] decoding these representations.
@@ 2.2 LLM Mechanism Interpretability
Jailbreak mechanism interpretability. Wei et al.[33] identify two primary factors behind jailbreak successes, i.e., competing objectives and mismatched generalization, in safety training. However, this work focuses predominantly on empirical findings, which treats the LLM as a black box.
@@
Several studies [15, 16, 17, 18] take pioneering steps in analyzing jailbreaks in the representation space. Lin et al. [15] propose that successful jailbreaks shift harmful prompts into clusters typically occupied by harmless ones. Ball et al.[16] examine similarities in activation vectors across different jailbreak attacks, suggesting that these commonalities reflect shared internal mechanisms for bypassing safety measures. Zhou et al. [18] find jailbreaks disrupt the association between harmfulness and negative emotion learned via safety alignment. Li et al.[17] identify that representations from middle layers of the model are crucial for distinguishing malicious queries from safe ones.
@@
Another type of studies [20, 19] analyzes the safety mechanism within LLMs at the level of circuit, particularly, the neurons in MLP layers. Chen et al.[20] propose activation contrasting to locate key neurons that are responsible for safety behaviors and dynamic activation patching to evaluate their causal effects. Zhao et al.[19] find a special neuron that changing the value of that neuron to 0 could impact the safe output of the model completely.
@@
Although representation analysis reveals shifts in latent space, it fails to identify the causes of these changes or pinpoint key components of vulnerability. While circuit analysis can locate specific vulnerable modules, it typically focuses on a limited set of components. However, jailbreaks often involve interactions across multiple layers, and their impact amplifies deeper into the model. Focusing on individual components makes it hard to capture the impact on both each layer and the model’s overall behavior.
To address these limitations, we conduct in-depth analysis of both representation and circuit perspectives and consider the diverse jailbreak strategies to understand better why and how these models are vulnerable towards jailbreak.
3 Preliminary

### v2

@@ Abstract.
Despite the outstanding performance of Large language Models (LLMs) in diverse tasks, they are vulnerable to jailbreak attacks, wherein adversarial prompts are crafted to bypass their security mechanisms and elicit unexpected responses. Although jailbreak attacks are prevalent, the understanding of their underlying mechanisms remains limited. Recent studies have explained typical jailbreaking behavior (e.g., the degree to which the model refuses to respond) of LLMs by analyzing representation shifts in their latent space caused by jailbreak prompts or identifying key neurons that contribute to the success of jailbreak attacks. However, these studies neither explore diverse jailbreak patterns nor provide a fine-grained explanation from the failure of circuit to the changes of representational, leaving significant gaps in uncovering the jailbreak mechanism. In this paper, we propose JailbreakLens, an interpretation framework that analyzes jailbreak mechanisms from both representation (which reveals how jailbreaks alter the model’s harmfulness perception) and circuit perspectives (which uncovers the causes of these deceptions by identifying key circuits contributing to the vulnerability), tracking their evolution throughout the entire response generation process. We then conduct an in-depth evaluation of jailbreak behavior on five mainstream LLMs under seven jailbreak strategies. Our evaluation reveals that jailbreak prompts amplify components that reinforce affirmative responses while suppressing those that produce refusal. This manipulation shifts model representations toward safe clusters to deceive the LLM, leading it to provide detailed responses instead of refusals. Notably, we find a strong and consistent correlation between representation deception and activation shift of key circuits across diverse jailbreak methods and multiple LLMs.
LLM Jailbreak, Interpretation, Representation, Circuit
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Security and privacy Software and application security
1. Introduction
@@
Large language models (LLMs) (Touvron et al., 2023; Yang et al., 2024) have revolutionized various fields with powerful capabilities. While LLMs acquire extensive knowledge from massive corpora during pre-training, they also inevitably embed harmful knowledge, which poisons LLMs to generate harmful responses that may violate ethics. To mitigate such harmful outputs, LLMs are further aligned via reinforcement learning to integrate security mechanisms (Ji et al., 2024; Bhardwaj et al., 2024; Huang et al., 2024) into them.
@@
However, flaws in the alignment process persist (Zou et al., 2023; Yu et al., 2023; Li et al., 2023b; Lv et al., 2024; Chao et al., 2023), as evidenced by numerous jailbreaks where attackers bypass safety mechanisms using carefully crafted prompts. Prevalent jailbreak methods can be categorized into five types (Jin et al., 2024), i.e., gradient-based attacks (Zou et al., 2023), evolutionary-based attacks (Yu et al., 2023; Liu et al., 2023), demonstration-based attacks (Li et al., 2023b), rule-based attacks (Lv et al., 2024; Ding et al., 2023), and multi-agent-based attacks (Chao et al., 2023). These attacks have raised concerns about LLM safety, for instance, in the widely reported cybertruck explosion outside the Trump Hotel in January 2025, investigations revealed that the perpetrator had acquired critical information by successfully jailbreaking ChatGPT (NewsHour, 2025).
@@
To understand the failure modes of safety alignment in real-world scenarios, interpreting the internal states of LLMs under jailbreak threats has become crucial for developing trustworthy superintelligence. Existing research (Lin et al., 2024; Ball et al., 2024; Li et al., 2024; Zhou et al., 2024) has explored how LLMs process and represent harmful prompts, showing that jailbreak prompts can be disguised within benign clusters, making it challenging for well-aligned models to distinguish between safe and harmful prompts. While representation analysis reveals the shifts in latent space, it cannot trace the specific internal components responsible for these changes. To address this, some studies (Zhao et al., 2023; Chen et al., 2024) focus on circuit analysis, particularly examining the neurons in MLP layers responsible for safety behaviors. Circuit analysis, on the other hand, usually focuses on a limited set of modules, such as certain neurons. Since jailbreak often involves interactions across multiple layers, the jailbreak effects are amplified deeper into the model, making it difficult to capture the impact on individual layers and overall behavior by analyzing only a limited set of circuits.
@@
Figure 1. An overview of JailbreakLens, a framework that interprets LLM jailbreak behavior via capturing representations and circuits. JailbreakLens first predicts the safety score of a specific representation and obtains two token sets that reflect affirmation and refusal in the representation space. We then use these token sets to attribute activation differences and identify key circuits. Finally, we combine the safety scores with key circuit activations to derive the final interpretation results.
While existing works have taken pioneering steps in exploring jailbreak mechanisms, we identify that there are still inherently several gaps. (1). Limitations in Jailbreak Types. Existing studies mainly focus on limited types of classic jailbreak methods, such as GCG and PAIR, etc., and thus fail to cover diverse state-of-the-art jailbreak strategies that exhibit novel attack features and limit the generability of the findings. (2). Insufficient Integration Between Representation and Circuits Analysis. Representation analysis reveals how jailbreaks deceive the model’s harmfulness perception, preventing it from recognizing malicious intent within the prompts, while circuit analysis uncovers the cause of this deception by identifying key circuits contributing to the vulnerability. However, existing words consider these two perspectives in isolation without combining two perspectives for more comprehensive interpretation. The ignored intergation obscures how high-level semantic deviations (representation-level) align with low-level functional failures (circuit-level). (3). Lack of Dynamic Analysis. Most studies focus on static analysis which examines the model behavior on the first token generation of jailbreak responses, failing to capture the fine-grained jailbreak behaviors throughout the sentence-level generation. Such lack leads to inaccurate interpretations. For example, a model may initially respond safely to refuse the malicious instruction, but gradually become affected by a jailbreak prompt and generate harmful content.
To address these gaps, we introduce JailbreakLens, overviewd in Fig. 1, a dual-perspective interpretability framework that systematically analyzes jailbreak behaviors from both the representation level and the circuit level. At the representation level, we examine how jailbreak prompts perturb harmfulness perception within the representation space with a safety probe. At the circuit level, we further localize and interpret the functional components responsible for safety behaviors. We extract the direction vector from the safety probes and project it into vocabulary space, obtaining two semantical token sets: affirmation tokens and refusal tokens. We then attribute the model’s internal activations to these token sets, aiming to identify key components that contributes to affirmation for harmless instructions and refusal for harmful instructions.
Our study reveals that jailbreak prompts succeed by subtly distorting the model’s internal understanding of harmfulness, allowing malicious instructions to be mistreated as safe within the representation space. This semantic misalignment is manipulated by a small set of safety-critical components that govern the model’s refusal or affirmation behaviors. Furthermore, we quantitatively demonstrate a strong and consistent correlation between representation deception and the activation patterns of key circuits, indicating that jailbreaks operate not through isolated perturbations, but through coordinated manipulation of both the model’s semantic perception and its functional circuits. Notably, our findings are generalizable across diverse jailbreak methods and multiple LLMs.
@@
  • Dual-Perspective Jailbreak Interpretation Method. We propose JailbreakLens, a novel interpretation framework that integrates representation-level and circuit-level analysis to interpret how jailbreak prompts distort harmfulness perception within representation space and disrupt safety-critical circuit activations, addressing the main gaps of jailbreak mechanism interpretation in current research.
  • Comprehensive Study of Jailbreak Strategies. We systematically evaluate seven jailbreak methods which covers the popular attack types on five mainstream LLMs, revealing consistent failure patterns that transcend attack types, model scales, and alignment strategies, demonstrating the generality of our interpretability approach.
  • Extensive Findings on Jailbreak Mechanism. Using our dual-perspective interpretation framework, we provide generalizable mechanism patterns across diverse jailbreak methods and multiple LLMs. We uncover that jailbreak prompts deceive models by shifting harmful inputs toward safe representation clusters, manipulating a small set of safety-critical components in deeper layers to suppress refusal behaviors and encourage affirmation, and also reveal a strong correlation between representation deception and key activations.
2. Related Works
@@ 2.1. Jailbreak Attacks on LLM
In jailbreak attcacks, attackers strategically manipulate the input prompts to bypass LLM safeguards. Existing jailbreak methods can be categorized into five main types based on the prompt generation strategy (Jin et al., 2024), i.e., gradient-based (Zou et al., 2023), evolutionary-based (Yu et al., 2023; Liu et al., 2023), demonstration-based (Li et al., 2023b), rule-based (Ding et al., 2023; Lv et al., 2024), and multi-agent-based (Chao et al., 2023) jailbreaks.
@@
Gradient-based jailbreaks exploit model gradients to craft input prompts. For example, GCG (Zou et al., 2023) optimizes a suffix that can elicit harmful responses when attached to a harmful query. The suffix is calculated by greedy search from random initialization to maximize the affirmative response likelihood.
@@
Evolutionary-based jailbreaks generate adversarial prompts via evolutionary strategies. For example, Yu et al. (Yu et al., 2023) propose GPTFuzzer, which initiate with human-crafted templates as seeds, and then mutate them to generate novel templates. Additionally, Liu et al. (Liu et al., 2023) introduce AutoDAN, which automatically generate stealthy jailbreak prompts using a hierarchical genetic algorithm.
@@
Demonstration-based jailbreaks use fixed prompt templates to elicit harmful responses from the target model. For example, DeepInception (Li et al., 2023b) leverages the personification capabilities of LLMs and induce the model to engage in complex, multi-layered contexts, subtly bypassing the safety guardrails.
@@
Rule-based jailbreaks decompose malicious prompts via predefined rules. For example, ReNeLLM (Ding et al., 2023) rewrites prompts while preserving core semantics and then nest them within universal task scenarios like table filling. CodeChameleon (Lv et al., 2024) circumvents safety guardrails by reformulating tasks into code completion format.
@@
Multi-agent-based jailbreaks depend on the cooperation of multiple LLMs to iteratively refine jailbreak prompts. For example, PAIR (Chao et al., 2023) use the collaboration of a red team LLM and an evaluator LLM to optimize the jailbreak prompts.
@@
To investigate diverse jailbreak strategies, we select 1-2 specific methods from each category, addressing the limited category exploration in previous work on jailbreak mechanism explanations.
2.2. LLM Mechanism Interpretability
Research on LLM mechanism interpretability aims to enhance the transparency of LLM behaviors through attribution methods, which are divided into input attribution (Deiseroth et al., 2023; Achtibat et al., 2024) and component attribution (Zhang and Nanda, 2023; Wang et al., 2022). Input attribution traces model responses to specific inputs, while component attribution (Zhang and Nanda, 2023; Wang et al., 2022) identifies influential model components via activation patching(Geiger et al., 2024; Hanna et al., 2024) or linear representation hypothesis(Park et al., 2023).
@@ 2.2. LLM Mechanism Interpretability
Jailbreak mechanism interpretability. Wei et al.(Wei et al., 2024) identify two primary factors for jailbreak success, i.e., competing objectives and mismatched generalization. However, they focus primarily on empirical findings, which treats LLM as a black box.
@@
Recent research on jailbreak mechanism interpretability has focused on representation-level (Lin et al., 2024; Ball et al., 2024; Li et al., 2024; Zhou et al., 2024) and circuit-level analysis. In representation-level analysis, researchers analyze jailbreaks within the model representation space. For example, Lin et al. (Lin et al., 2024) propose that successful jailbreaks shift harmful prompts into harmless clusters. Ball et al.(Ball et al., 2024) demonstrate that different jailbreak methods share similiar activation vectors for bypassing safety guardrails. Zhou et al. (Zhou et al., 2024) find jailbreaks disrupt the association between harmfulness and negative emotion. Li et al.(Li et al., 2024) identify that the representation in middle layers is crucial for distinguishing malicious queries from safe ones.
@@
Circuit-level analysis examine LLM safety mechanisms at the circuit level, particularly, the neurons in MLP layers. Chen et al.(Chen et al., 2024) propose activation contrasting to locate key neurons that are responsible for safety behaviors. Zhao et al.(Zhao et al., 2023) find a critial neuron whose value significantly impacts the safe output of the model.
@@
Although representation analysis reveals shifts in latent space, it fails to identify the causes of these changes or pinpoint key components of vulnerability. While circuit analysis can locate specific vulnerable modules, it typically focuses on a limited set of components despite jailbreaks involving multi-layer interactions, making it hard to capture the impact on both each layer and the model’s overall behavior. To address these limitations, we conduct in-depth analysis of both representation and circuit perspectives and consider the diverse jailbreak strategies to understand better why and how these models are vulnerable towards jailbreak.
3. Preliminary